It is anticipated that the temperature in the construction site changes from 0 oC to 33 oC. If one of the
components to be installed in the construction sites is a steel member of 600 meters in length, how
much expansion should be anticipated?
1
Expert's answer
2020-11-27T10:14:37-0500

We can find the change in the linear expansion from the formula:


"\Delta L=\alpha L\Delta T,""\Delta L=1.2\cdot10^{-5}\ ^{\circ}C^{-1} \cdot 600\ m\cdot 33\ ^{\circ}C=0.24\ m."

Answer:

"\Delta L=0.24\ m."